OpenStudy (anonymous):

If y = 8 cm and z = 15 cm, what is the length of x?

OpenStudy (kdr1308):

Was there a picture?

OpenStudy (anonymous):

Yes

OpenStudy (kdr1308):

Thanks

OpenStudy (anonymous):

Your welcome did that go through?

OpenStudy (kdr1308):

Yes!

OpenStudy (anonymous):

Okayy

OpenStudy (abdullahm):

Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

Yes isnt it like A2+b2=c2?

OpenStudy (abdullahm):

Correct! But for your question, it would be: \(\bf z^2 + y^2 = x^2\)

OpenStudy (abdullahm):

So, simply plug in z = 15, and y = 8, and solve for x!

OpenStudy (anonymous):

Okay!

OpenStudy (abdullahm):

Tell me what you get so I can make sure your answer is correct! :)

OpenStudy (anonymous):

I got 17

OpenStudy (abdullahm):

That is correct! :)

OpenStudy (anonymous):

Yay! Thanks so much! <3

OpenStudy (abdullahm):

It was my pleasure! :)

OpenStudy (ace-n-it):

good job